OpenStudy (anonymous):

Why isn't 1 a prime number? :)

OpenStudy (alexwee123):

because there is only one factor for 1 and a prime number requires 2

OpenStudy (anonymous):

caz many theorem will fail

jimthompson5910 (jim_thompson5910):

The fundamental theorem of arithmetic states that you can represent any number uniquely as a product of primes If 1 were prime, then a number like 12 can be written as 12 = 2*2*3 12 = 1*2*2*3 12 = 1*1*2*2*3 ...etc so 12 is being represented as a product of primes but it's not unique. So make the representation unique, we exclude 1 from being prime Note: the order of the factors doesn't matter

OpenStudy (anonymous):

Beacause its too COOL

OpenStudy (anonymous):

because nothing can go into it besides itself

OpenStudy (anonymous):

Every prime number has exactly two factors! For example, 2 is prime because it only has factors 1 and 2. 3 only has factors 1 and 3. 5 only has factors 1 and 5. Now, most non-primes, it's easy to realize, "Oh, okay, it has more than 2 factors." 1 is the unique case where it has fewer than 2 factors. Just 1 =/

OpenStudy (anonymous):

THank you (: